# Advancing Mobile GUI Agents: A Verifier-Driven Approach to Practical Deployment

### v2

@@ 4.1. Decision-Making Training
Training loss design. The following training loss is adopted to train the generative verifier at each step in V-Droid:
@@
where $\alpha_{\kappa}^{n}(t)\in\mathcal{A}^{\prime}_{\kappa}(t)$ and $\tau$ represents the assigned score estimation for the particular verified action. Specifically, in V-Droid, a trainable MLP layer is appended to the LLM to project its output token probabilities into the estimated score for each action.
@@
When the $P^{3}$ loss $L_{\kappa}(t)$ at each step $t$ of task $\kappa$ across all labeled tasks in the training dataset is minimized during training, the verifier learns to maximize the score assigned to the positive action while minimizing the score assigned to the negative action.
In P3 training, we introduce three key concepts: pairwise, process, and preference, all of which are crucial for effectively training the verifier. Firstly, preference training offers greater discriminative ability for the verifier to distinguish similar UI or actions which can be otherwise misleading to mobile agents. Secondly, process supervision provides fine-grained supervision signals, which are inherently beneficial for improving the step-wise decision-making ability of the verifier. Furthermore, compared to outcome-based supervision , process supervision allows for the construction of a larger number of training samples. For instance, from 100 task traces, we can derive around 55K step-wise process labels. Thirdly, pairwise training can be considered a form of data augmentation. Without such augmentation, only one training sample might be obtained for each task step. In contrast, pairwise training enables the generation of 50$\times$ more training samples, which is especially critical when the available training data is limited.
@@ 4.2. Self-Correcting Training
Such data are valuable in enhancing the self-correcting capability of mobile agents. Intuitively, when humans interact with devices, it is entirely possible to enter an erroneous state unrelated to the current intent. Instead of abandoning the task and starting over, humans often try to correct such mistakes, e.g., by clicking the navigate back button. To the best of our knowledge, no existing mobile agent possesses the self-correcting capability. But the action pairs derived from these erroneous states present an opportunity for us to equip V-Droid with self-correcting ability through $P^{3}$ training.
Specifically, as illustrated in Fig. 6, given that the state $\mathcal{S}_{t-1}$ at step $t-1$ lies on the correct trace, the state $\mathcal{S}_{t}$ can be reached by executing the labeled correct action $\alpha_{\kappa}^{y}(t-1)$. Conversely, executing any incorrect action $\alpha_{\kappa}^{n}(t-1)\in\mathcal{A}_{\kappa}^{\prime}(t-1)$ would lead to an erroneous state for the given task $\kappa$, noted as $\mathcal{S}^{\prime}(t)$.
On the erroneous state $\mathcal{S}^{\prime}(t)$, none of the actions from the action space may be applicable. Instead, a reverse action $\alpha_{\kappa}^{r}(t)$ is introduced as the correct action, e.g., pressing the navigate back button. Specifically, the reverse action is derived based on the executed action at the previous step $t-1$ through a predefined mapping relationship, as illustrated in Fig. 6. For example, when V-Droid enters incorrect content by executing the action type text, the corresponding reverse action is to clear the previously input content.
@@
(5) $\Gamma^{\prime}_{\kappa}(t)=\{(\rho_{\kappa}^{t}(\alpha_{\kappa}^{r}(t)),\rho_{\kappa}^{t}(\alpha_{\kappa}(t)))\;|\;\alpha_{\kappa}(t)\in\mathcal{A}_{\kappa}(t)\},$
where $\alpha_{\kappa}^{r}(t)$ represents the derived reverse action, particularly
@@ 5. Human-Agent-Joint-Annotation
$P^{3}$ training significantly enhances the capabilities of mobile agents; however, the necessary training data—particularly fine-grained process labels that capture both correct states and erroneous states—remains, to the best of our knowledge, unavailable in public repositories. To this end, we propose a novel data synthesis methodology, as illusrated in Fig. 7. First, we collect task instructions for Android applications from public datasets (Li et al., 2024a; Rawles et al., 2023). Additionally, we leverage LLMs to synthesize additional task instructions based on the applications’ descriptions available on the App Store, followed by meticulous manual verification. Subsequently, we generate the training data by actually executing the targeted tasks using the corresponding applications on our culster of Android emulators. Human and agents jointly annotate the correct actions required at each step towards successful task completion.
@@ 5.2. Iterative Annotation and Training
Leveraging the human-agent joint annotation scheme, we iteratively optimize the verifier using $P^{3}$ training, deploying it to collect new data for the next round on out-of-domain tasks and APPs.
V-Droid undergoes four rounds of data collection and training. To address the cold start problem in the first round, we manually label data without agent assistance. Human annotators are instructed to select the correct actions from the extracted action lists provided by V-Droid. In the subsequent rounds, the agent is deployed to execute tasks and mark step-wise correctness, progressively scaling the dataset size from 9K to 27K, then to 55K and 110K at the end of the 4-th iterations.
As data volume grows, the verifier demonstrates continuous performance improvements in both decision-making capability and annotation accuracy. Table 1 presents the truth table across different iterations, comparing verifier predictions based on entropy with step-wise ground truth. The true positive ratio increases from 0.11 to approximately 0.40, while annotation accuracy improves from 0.51 to over 0.70.
These improvements indicate that as the verifier’s scoring ability strengthens, the agent correctly annotates more steps, further reducing human annotation effort and enhancing training efficiency.
Table 1. Truth table based on entropies of scores on different iterations. The threshold is set as the median value of the entropies.
@@ 5.3. Training Details
In V-Droid, the generative verifier is built based on Llama-3.1-8B-Instruct-4-bit with one MLP layer attached to project the token logits into the action scores. We use Q-LoRa training with rank 16. The learning rate is set to 1e-4 with 10 warm-up steps and then gradually reduced to 1e-6 until the end of the training. The training epoch is set to 20 for the 4th iteration, which is larger than the usual LoRa training epoch. The reason is that longer training enlarge the score gap between the correct and rejected actions, which is observed to be more robust during the test time. The 4th iteration of training is conducted on around 110k data pairs for 90 hours on 16×16$\times$16 ×Nvidia A100 40G GPUs.
@@
(a) Decision-making latency
(b) Step-wise latency
Figure 9. Decision-making latency and the total latency per step of V-Droid compared to typical mobile agents.
@@ 6.1.1. Benchmark
MobileAgentBench (Wang et al., 2024a) provides a realistic mobile phone usage environment built upon 10 open-source applications and 100 tasks.
In contrast to static benchmarks, e.g., DroidTask (Wen et al., 2024a) and AndroidControl (Li et al., 2024a), the realistic and dynamic Android environment we evaluate V-Droid can better reflect its ability. Among the three benchmarks, AndroidLab and MobileAgentBench are designated as out-of-domain (OOD) test sets. During the training data collection process, we rigorously excluded applications utilized in these two benchmarks to ensure unbiased evaluation.
@@ 6.1.2. Baselines
V-Droid is compared with a wide range of mainstream mobile agents, including text-only agents (T3A (Rawles et al., 2024), AutoDroid (Wen et al., 2024a), Ponder&Press (Wang et al., 2024e)), multimodal agents (M3A (Rawles et al., 2024), SeeAct (Zheng et al., 2024), AndroidArena (Xing et al., 2024), CogAgent (Hong et al., 2024), APPAgent (Li et al., 2024b), MobileAgent (Wang et al., 2024d)), and agents with grounding models (Agent-S2 (Agashe et al., 2025), Aria-UI (Yang et al., 2024), UGround (Gou et al., 2025), and UI-TARS (Qin et al., 2025)). Advanced LLMs, including GPT-3.5, GPT-4, GPT-4V, GPT-4o, GLM, DeepSeek-R1, Qwen, Llama-3.1, Claude and Gemini, are included as the base models for these agents for the comparison.
@@ 6.1.3. Metrics
V-Droid is evaluated against the baselines using two key metrics: task success rate (SR) and latency. For the SR of the baselines, we directly adopt the SR number reported in their respective papers on the corresponding benchmarks. To assess latency, we employ two measures: the decision-making latency of the LLM and the total step-wise latency of the entire agent. The total step-wise latency primarily comprises the decision-making latency and the working memory update latency. Latency is evaluated using 20 randomly sampled tasks from benchmarks.
@@ 6.1.4. Evaluation platform
We evaluate V-Droid across various hardware configurations. The agent system runs within an Android emulator on a PC equipped with an Intel i9-10900X CPU. The LLMs used by V-Droid and baseline models are tested on NVIDIA GPUs, including 4090, A100, A6000. Unless otherwise specified, all reported time measurements of V-Droid are conducted on a server with two NVIDIA 4090 GPUs.
@@ 6.2. Performance Improvement
Improved task success rate. Fig. 8 demonstrates that V-Droid outperforms existing mobile agents across three realistic MobileAgentBenchmarks, AndroidWorld, AndroidLab, and MobileAgentBench, achieving success rate (SR) improvements of 9.5%, 2.1%, and 9.0%, respectively. Compared to cloud-based LLM-powered agents (e.g., GPT-4, GPT-4o, DeepSeek-R1, Gemini-1.5-Pro, Claude-3.5), V-Droid achieves 25.0% and 7.13% higher SR on AndroidWorld and AndroidLab, respectively. Against advanced mobile agent frameworks (e.g., Agent-S2, AutoDroid, AppAgent, SeeAct, CogAgent), V-Droid improves SR by 9.5% on AndroidWorld and 9.0% on the MobileAgentBench. Compared to models that decompose decision-making into reasoning and grounding (e.g., UI-TARS, Aria-UI, UGround, Aguvis), V-Droid demonstrates a 14.3% SR improvement on the AndroidWorld. Against other fine-tuned SLMs (e.g., Qwen-VL-7B-FT, Llama-3.1-8B-FT), V-Droid achieves a notable 14.3% improvement on AndroidLab. Unlike existing generation-based GUI agents that operate in continuous UI spaces, V-Droid simplifies decision-making by mapping UI states to a finite action space and separating the process into verification and completion, making it more manageable for SLMs. Furthermore, the $P^{3}$ training framework, which utilizes large-scale data, improves the verifier’s ability to distinguish similar actions, self-correct errors, and maintain awareness of progress.
@@
Reduced Latency. Fig. 9 highlights the significant speed advantage of V-Droid over SOTA mobile agents in both decision-making and step-wise latency. While SOTA agents typically take over 20 seconds per step, V-Droid averages just 0.7 seconds to generate an action based on UI descriptions and working memory, and 3.8 seconds to complete a full step. Notably, in 88% of cases, where selected actions require no further completion, decision-making takes only 0.44 seconds. Compared to grounding-based agents that decompose decision-making into reasoning and grounding (e.g., UI-TARS, Aria-UI, UGround, Aguvis), V-Droid achieves up to a 32.1×32.1$\times$32.1 × speedup. Against UI-TARS (Qin et al., 2025) and DeepSeek-R1 with System-2 reasoning (Guo et al., 2025), V-Droid achieves 25.1×25.1$\times$25.1 × and 112.2×112.2$\times$112.2 × speed improvements, respectively. This efficiency gain stems from V-Droid ’s verifier-driven workflow, which transforms traditional auto-regressive decoding into a parallelized, prefilling-only scoring process. Additionally, verifier prompts are optimized to share most of their content across different actions, allowing KV caching for the shared prefix, significantly accelerating scoring speed. To the best of our knowledge, V-Droid is the first mobile GUI agent capable of near-real-time decision-making. While the decision-making speed is fast, its overall per-step latency is primarily constrained by working memory construction, averaging 3.03 seconds, where UI description extraction and action retrieval take less than 0.5 ms. We further discuss design alternatives for optimizing working memory in Section 6.4.
Self-correction showcase. As shown in Fig. 10, V-Droid explores one reasonable action ”Click Textfile.txt” but then realizes itself in a wrong status that deviates from the goal. Later, it selects to navigate back and long-press the button to reveal the file properties. Similar cases are observed on the other two benchmarks, which can help improve the success rate of V-Droid by 7.3% totally.
@@
Figure 10. A showcase of V-Droid automating the task ’Check the file property of Textfile.txt’ on MobileAgentBench. Note that both this task and the corresponding application were excluded from our training process.

### v3

@@ 4.1. Decision-Making Training
Training loss design. The following loss is adopted to train the generative verifier at each step in V-Droid:
@@
where $\alpha_{\kappa}^{n}(t)\in\mathcal{A}^{\prime}_{\kappa}(t)$ and $\tau$ represents the assigned score for the verified action. Specifically, in V-Droid, a trainable MLP layer is appended to the LLM to project its output token probabilities into the estimated score for each action.
@@
When the $P^{3}$ loss $L_{\kappa}(t)$ at each step $t$ of task $\kappa$ across all labeled tasks is minimized during training, the verifier learns to maximize the score assigned to the positive action while minimizing the score assigned to the negative action.
[…]
@@ 4.2. Self-Correcting Training
Such data are valuable in enhancing the self-correcting capability of mobile agents. Intuitively, when humans interact with devices, it is possible to enter an erroneous state unrelated to the current intent. Instead of abandoning the task and starting over, humans often try to correct such mistakes, e.g., by clicking the navigate back button. To the best of our knowledge, no existing mobile agent possesses the self-correcting capability. But the action pairs derived from these erroneous states present an opportunity for us to equip V-Droid with self-correcting ability through $P^{3}$ training.
Specifically, as illustrated in Fig. 6, given that the state $\mathcal{S}_{t-1}$ at step t1t-1 lies on the correct trace, the state $\mathcal{S}_{t}$ can be reached by executing the labeled correct action $\alpha_{\kappa}^{y}(t-1)$. Conversely, executing any incorrect action $\alpha_{\kappa}^{n}(t-1)\in\mathcal{A}_{\kappa}^{\prime}(t-1)$ would lead to an erroneous state for the given task $\kappa$, noted as $\mathcal{S}^{\prime}(t)$.
On the erroneous state $\mathcal{S}^{\prime}(t)$, none of the actions from the action space may be applicable. Instead, a reverse action $\alpha_{\kappa}^{r}(t)$ is introduced as the correct action, e.g., pressing the navigate back button. Specifically, the reverse action is derived based on the executed action at the previous step t1t-1 through a predefined mapping relationship, as illustrated in Fig. 6. For example, when V-Droid enters incorrect content by executing the action type text, the corresponding reverse action is to clear the previously input content.
@@
(5) $\Gamma^{\prime}_{\kappa}(t)=\{(\rho_{\kappa}^{t}(\alpha_{\kappa}^{r}(t)),\rho_{\kappa}^{t}(\alpha_{\kappa}(t)))\;|\;\alpha_{\kappa}(t)\in\mathcal{A}_{\kappa}(t)\}.$
The $\alpha_{\kappa}^{r}(t)$ represents the derived reverse action,
@@ 5. Human-Agent-Joint-Annotation
$P^{3}$ training enhances the capabilities of mobile agents. However, to the best of our knowledge, the necessary training data, particularly fine-grained process labels that capture both correct and erroneous states, remains unavailable in public repositories. To this end, we propose a novel data synthesis methodology, as illusrated in Fig. 7. First, we collect task instructions for Android Apps from public datasets (Li et al., 2024a; Rawles et al., 2023). Additionally, we leverage LLMs to synthesize additional task instructions based on the Apps’ descriptions available on the App Store, followed by meticulous manual verification. Subsequently, we generate the training data by actually executing the targeted tasks using the corresponding applications on our cluster of Android emulators. Human and agents jointly annotate the correct actions required at each step towards successful task completion.
@@ 5.2. Iterative Annotation and Training
(a) Distribution of training data across app. categories
(b) Distribution of available actions per step
Figure 8. Statistics of the collected datasets for training.
Leveraging human–agent joint annotation, we iteratively optimize the verifier using $P^{3}$ training and deploy it to collect new data on out-of-domain tasks and apps over four rounds. To address the cold-start problem in the first round, we manually labeled data without agent assistance. Human annotators selected the correct actions from the extracted action lists provided by V-Droid, requiring approximately four minutes per task and about five hours in total to collect the initial 9K data pairs. In subsequent rounds, the agent was deployed to execute tasks and mark stepwise correctness, enabling progressive dataset expansion from 9K to 27K, then to 55K, and finally to 110K after four iterations. Human annotators primarily corrected errors in the trajectories, which reduced the average annotation time to around one minute per task. As shown in Fig. 8, the collected dataset spans more than 90 applications across seven common categories.
As data volume grows, the verifier demonstrates continuous performance improvements in both decision-making capability and annotation accuracy. Table 1 presents the truth table across different iterations, comparing verifier predictions based on entropy with step-wise ground truth. The true positive ratio increases from 0.11 to approximately 0.40, while annotation accuracy improves from 0.51 to over 0.70. These improvements indicate that as the verifier’s scoring ability strengthens, the agent correctly annotates more steps, further reducing human annotation effort and enhancing training efficiency.
Table 1. Truth table based on entropies of scores. The threshold is set as the median value of the entropies.
@@ 5.3. Training Details
In V-Droid, the generative verifier is built based on Llama-3.1-8B-Instruct-4-bit with one MLP layer attached to project the token logits into the action scores. We use Q-LoRa training with rank 16. The learning rate is set to 1e-4 with 10 warm-up steps and then gradually reduced to 1e-6 until the end of the training. The training epoch is set to 20 for the 4th iteration, which is larger than the usual LoRa training epoch. The reason is that longer training enlarge the score gap between the correct and rejected actions, which is observed to be more robust during the test time. The 4th iteration of training is conducted on around 110k data pairs for 90 hours on $16\timesNvidia$ A100 40G GPUs.
@@
(a) Step-wise latency
(b) Decision-making latency
Figure 10. The step-wise latency and decision-making latency of V-Droid compared to typical mobile agents.
@@ 6.1.1. Benchmark
MobileAgentBench (Wang et al., 2024a) provides a phone usage environment built on 10 open-source applications and 100 tasks.
In contrast to static benchmarks, e.g., DroidTask (Wen et al., 2024a) and AndroidControl (Li et al., 2024a), the realistic and dynamic Android environment we evaluate V-Droid can better reflect its ability. Among the three benchmarks, AndroidWorld is designated as the in-the-domain test set, whereas AndroidLab and MobileAgentBench are designated as out-of-domain (OOD) test sets. During the training data collection process, we rigorously excluded applications utilized in these two benchmarks to ensure unbiased evaluation. The differences in data distributions between the training set and the test benchmarks are illustrated in Fig. 8(b).
@@ 6.1.2. Baselines
V-Droid is compared with a wide range of mainstream mobile agents, including text-only agents (T3A (Rawles et al., 2024), AutoDroid (Wen et al., 2024a), AutoDroid-V2 (Wen et al., 2024b), MobileGPT (Lee et al., 2024), Ponder&Press (Wang et al., 2024e)), multimodal agents (M3A (Rawles et al., 2024), SeeAct (Zheng et al., 2024), AndroidArena (Xing et al., 2024), CogAgent (Hong et al., 2024), APPAgent (Li et al., 2024b), MobileAgent (Wang et al., 2024d)), and agents with grounding models (Agent-S2 (Agashe et al., 2025), Aria-UI (Yang et al., 2024), UGround (Gou et al., 2025), and UI-TARS (Qin et al., 2025)). Advanced LLMs, including GPT-3.5, GPT-4, GPT-4V, GPT-4o, GLM, DeepSeek-R1, Qwen, Llama-3.1, Claude and Gemini, are included as the base models for these agents.
@@ 6.1.3. Metrics
V-Droid is evaluated against baselines using two key metrics: task success rate (SR) and latency. We directly adopt the SR number reported in the respective baseline papers on the corresponding benchmarks. To assess latency, we employ two measures: the total step-wise latency of the entire agent and the decision-making latency of the LLM. The total step-wise latency primarily comprises the decision-making latency, the working memory update latency, and the execution time. Latency is evaluated using 20 randomly sampled tasks from benchmarks.
@@ 6.1.4. Evaluation platform
We evaluate V-Droid across various hardware configurations. The agent system runs within an Android emulator on a PC with an Intel i9-10900X CPU. The LLMs used by V-Droid and baseline models are tested on NVIDIA GPUs, including 4090, A100, A6000. Unless otherwise specified, all time measurements of V-Droid are conducted on a server with two NVIDIA 4090 GPUs.
@@ 6.2. Performance Improvement
Improved task success rate. Fig. 9 demonstrates that V-Droid outperforms existing mobile agents across three realistic MobileAgentBenchmarks, AndroidWorld, AndroidLab, and MobileAgentBench, achieving success rate (SR) improvements of 5.2%, 2.1%, and 9.0%, respectively. Compared to cloud-based LLM-powered agents (e.g., GPT-4, GPT-4o, DeepSeek-R1, Gemini-1.5-Pro, Claude-3.5), V-Droid achieves 25.0% and 7.13% higher SR on AndroidWorld and AndroidLab, respectively. Against advanced mobile agent frameworks (e.g., Agent-S2, AutoDroid, AppAgent), V-Droid improves SR by 5.2% on AndroidWorld and 9.0% on the MobileAgentBench. Compared to models that decompose decision-making into reasoning and grounding (e.g., UI-TARS, Aria-UI, UGround, Aguvis), V-Droid demonstrates a 14.3% SR improvement on the AndroidWorld. Against other fine-tuned SLMs (e.g., Qwen-VL-7B-FT, Llama-3.1-8B-FT), V-Droid achieves a notable 14.3% improvement on AndroidLab. Compared with memory-driven agents like AutoDroid and MobileGPT, V-Droid achieves over 36.2%, 21.0%, and 18.0% SR improvement on the three benchmarks.
Unlike existing generation-based GUI agents that operate in continuous UI spaces, V-Droid simplifies decision-making by mapping UI states to a finite action space and decomposing the process into verification and completion, making it more tractable for SLMs. In addition, the verifier-driven workflow embeds UI-specific knowledge directly into the action candidates, thereby reducing the complexity of decision-making. Moreover, the $P^{3}$ training framework, built on large-scale data, further enhances the verifier’s ability to distinguish between similar actions, self-correct errors, and maintain awareness of task progress.
Table 2. Decision-making latency of GPT-4o, UI-TARS-7B (Qin et al., 2025) with and without CoT, compared to V-Droid.
Agent GPT-4o UI-TARS-7B V-Droid
Decision w/ CoT w/o CoT w/ CoT w/o CoT Verifier
SR (%) 41.3 39.2 33.2 29.3 59.5
Latency (s) 6.70 5.74 10.6 0.97 0.74
Input Tokens 6.2K 6.2K 8.9K 2.9K 2.6K
Output Tokens 63.4 19.6 75.9 14.6 54.3
Reduced Latency. Fig. 10 highlights the significant speed advantage of V-Droid over SOTA mobile agents in both step-wise and decision-making latency. While SOTA agents typically take over 20 seconds per step, V-Droid takes just 4.3 seconds per step, which is $6.1\times$ faster. Specifically, it takes 0.44s for the verification (stage ③ in Fig. 4), 0.30s for the action completion (stage ④), and 3.03s for the working memory (stage ⑤). The action space and the prompt construction takes less than 1ms (stage ①-②) and the photo transition time takes 0.54s on average.
@@
This efficiency gain stems from V-Droid ’s verifier-driven workflow for decision-making, which transforms traditional auto-regressive decoding into a parallelized, prefilling-only scoring process. As shown in Fig. 10(b), V-Droid achieves up to a $32.1\times$ speedup compared to grounding-based agents that decompose decision-making into reasoning and grounding (e.g., UI-TARS, Aria-UI, UGround, Aguvis). Against UI-TARS (Qin et al., 2025) and DeepSeek-R1 with System-2 reasoning (Guo et al., 2025), V-Droid achieves $25.1\times$ and $112.2\times$ speed improvements, respectively. To the best of our knowledge, V-Droid is the first mobile GUI agent capable of near-real-time decision-making. We further discuss design alternatives for optimizing working memory in Section 6.4.
@@
Figure 11. A showcase of V-Droid automating the task ’Check the file property of Textfile.txt’ on MobileAgentBench. Note that both this task and application were excluded from the training process
The action completion in V-Droid is only needed in 12.4% cases. In some input heavy tasks, the stepwise latency of V-Droid could be longer. For instance, in calendar-related tasks, there are 25.5% actions needs completion, which increases the average latency to 0.71s for the completion and 4.7s per step. In contrast, selection-heavy tasks that do not require text input (such as opening Wi-Fi) achieve 4.0s per step.
Comparison with CoT-Disabled Agents. The advantages of the verification-driven workflow in V-Droid are further demonstrated in Table 2 on AndroidWorld. Disabling CoT reasoning in GPT-4o and UI-TARS leads to reduced decision-making latency. However, this comes at the cost of decreased SR. Despite removing CoT, both GPT-4o and UI-TARS still exhibit higher latency than V-Droid, as they autoregressively generate full actions. In contrast, V-Droid achieves a substantially higher SR with lower decision-making latency. These results further underscore the generation– verification gap in GUI agents and highlight the efficiency and effectiveness of verifier-driven decision-making.
Self-correction showcase. In Fig. 11, V-Droid explores one reasonable action ”Click Textfile.txt” but then realizes itself in a wrong status that deviates from the goal. Later, it selects to navigate back and long-press the button to reveal the file properties. Similar cases are observed on the other two benchmarks. Notably, the self-correction training improves the SR of V-Droid from 52.2% to 59.5% on AndroidWorld and slightly increases the average trajectory length from 10.3 to 11.6 steps due to additional explorations. This modest increase in trajectory length is a reasonable trade-off given the substantial gain in overall success rate.